OpenStudy (pphalke):

If a right triangle has one leg that measures 12 centimeters, and a hypotenuse that measures 13 centimeters, how long is the other side?

OpenStudy (lukesolo_2187):

a^2+b^2=c^2 12^2+b^2=13^2 144+b^2=169 144-144+b^2=169-144 b^2=25 square root of b^2=square root of 25 b=5

OpenStudy (lukesolo_2187):

answer is 5.

OpenStudy (lukesolo_2187):

can i get medal or something

OpenStudy (lukesolo_2187):

thanks